The solution of s/3-5/6=s+2/9 is?
Answer by Alan3354(69443)

The solution of s/3-5/6=s+2/9 is?
-------------
Solve for s?
Multiply by 18 to get rid of fractions
6s - 15 = 18s + 4
-12s = 19
s = -19/12